# Synchronizing microwave cQED limit-cycle oscillators

### v1

@@ II.2 Effective photon action to fourth order in $g$
Figure 2: Real and imaginary parts of the self energy plotted in panels (a, c) against DQD detuning for $V=10\omega_{0}$, and in panels (b, d) against bias voltage for $\varepsilon=(\sqrt{3}/2)\omega_{0}$. In both cases our $L/R$-symmetric parameter set implies that $\Pi^{\text{R}}_{LL}=\Pi^{\text{R}}_{RR}$ and $\Pi^{\text{R}}_{LR}=\Pi^{\text{R}}_{RL}.$ The dashed gridline in (c) shows the resonance condition, $\varepsilon=(\sqrt{3}/2)\omega_{0}$, and in (d) the threshold condition for the bias voltage, $V=\omega_{0}.$ The width of the transition region around the dashed gridline is set by $\Gamma$.
@@ III.2 Fluctuations: Fokker-Planck and beyond
Figure 4: Limit cycle radius, $r_{\ast}$, predicted respectively by the saddle-point equations, $r_{\text{sp}}$, or the value of Wigner function maximum, $r_{\text{W}}$, and plotted together with the root mean square of $r$ calculated using the LME. The inset shows a zoom in on the bifurcation region, $g\sim g_{c}$, revealing a slight difference in $g_{c}$ between the LME, and the vanishinig of $\tilde{\kappa}$ (marked by gridline).
@@
for a distribution function, $P(\bar{\phi},\phi,t)$ and with diffusion constant $D_{0}$. Transforming to polar representation and employing Wirtinger derivatives, $\partial_{\phi}=e^{-i\theta}(\partial_{r}-ir^{-1}\partial_{\theta})/2$ and $\partial_{\bar{\phi}}=e^{i\theta}(\partial_{r}+ir^{-1}\partial_{\theta})/2$, this transforms to
@@
where $c$ ensures normalization, $\int_{0}^{\infty}\!dr\,2\pi rP_{0}(r)=1$, and the subscript 0 indicates that $\Lambda_{2,3,4,5}$ are all omitted. The same solution was found in Ref. [12], for a LME with jump operators representing 1-photon drive and 2-photon losses. Without the nonlinearity, one can relate the distribution function to the Wigner function as $P(\bar{\phi},\phi,t)=W(\bar{\phi}/\sqrt{2},\phi/\sqrt{2},t)/2$. In the general nonlinear case, however, this identification is not valid. In particular, we note that the solution (27) is not a true Wigner function in the sense that the corresponding density matrix is not strictly positive semi-definite.
The solution (27) is valid only if $\Lambda_{1}^{\prime\prime}/D_{0}>0$, which is the case for $V>\omega_{0}$ (cf. Fig. 3), in which case the Hopf bifurcation takes place as $\tilde{\kappa}$ passes through zero. Below and above the bifurcation, the maximum of the probability distribution is attained, respectively, at $r=0$, or $r=r_{\ast}=\sqrt{-\tilde{\kappa}/\Lambda_{1}^{\prime\prime}}$, consistent with the saddle point solution found above. This bifurcation is shown in Fig. 4 as a sharp transition from the trivial fixed point, $r=0$, to the limit cycle at $r=r_{\ast}$ at a critical coupling, $g_{c}$, close to and above which $r_{\ast}\sim|g-g_{c}|^{1/2}$. The orange curve is determined as the fixed point of the classical saddle point equations, corresponding to the $r$ at which $P_{0}(r)$ reaches its maximum. The dark blue curve shows the exact radius of the limit cycle obtained as the maximum of the Wigner function calculated using the LME, which treats the interaction between photons and electrons exactly. From the inset, the true bifurcation is seen to take place for a slightly smaller coupling than the value $g_{c}=0.043\omega_{0}$, determined from the saddle-point equations by the vanishing of $\tilde{\kappa}=0$. Fig. 4 also shows that the saddle-point value for the limit-cycle radius, $r_{\text{sp}}$, quickly becomes inaccurate for $g/\omega_{0}\gtrsim 0.06.$
@@
We solve this eigenvalue problem numerically  [^1] The calculations were performed using the Mathematica function NDeigensystem, where the area of the cells in the discrete spatial grid can be specified., and find that except for the zero eigenvalue corresponding to the steady state, all eigenvalues of the resulting differential operator $\mathcal{L}_{1}+\mathcal{L}_{2}$ are real and negative. Fig. 5 shows a comparison of the steady state solution, $P$, to the previous solution, $P_{0}$, obtained from Eq. (27), and to the exact Wigner function, $W$, obtained from the LME. While all solutions preserve the bifurcation, the inclusion of $\Lambda_{2}$ and $\Lambda_{5}$ clearly improves the approximate solution and is seen (cf. panel d) to lower $g_{c}$ towards the correct value. Note that for the parameters considered here, the linear shift performed earlier, $\delta X^{\rm cl}=2gn_{R}/\omega_{0}$, is of the order of $10^{-2}$ and therefore insignificant in these plots.
Figure 5: Wigner function [(a), (c), (d)] and radial probability distribution [(b), (d), (f)] obtained by solving the FP equation without/with ($P_{0}/P$) the nonlinear terms ($\Lambda_{2/5}$), plotted alongside the Wigner function ($W$). From top to bottom the coupling strength increases as $g/\omega_{0}=[0.01,0.042,0.05]$. The dashed gridline in panel (f) marks the maximum of $P_{0}$, i.e. the saddle-point value of the limit-cycle radius.
@@ III.3 Entrainment
we solve for $(X(t),P(t))=({\rm{Re}}[\phi],{\rm{Im}}[\phi])/\sqrt{2}$. The corresponding phase-space trajectories are shown in Fig.6 together with the Lissajous curves traced out by $(X(t),F(t))$, revealing frequency entrainment when the amplitude of the force exceeds a critical strength, found here to be $A_{c}=0.044$. The rightmost column shows the power spectral density (PSD), calculated from the auto-correlation function of $X(t)$ (cf. Appendix B), which displays a clear transfer of spectral weight from $\tilde{\omega}_{0}$ to $\Omega$ as the amplitude of the force is increased. Note that the entrainment can also be found below the bifurcation point, $g<g_{c}.$ In this case any finite driving amplitude makes it possible to sustain a closed phase-space trajectory with finite radius, despite the fixed point of the undriven system being $r=0$ (see Fig. 6(d)). The radius of the trajectory increases with increasing $A,$ and decreasing detuning, $|\Omega-\tilde{\omega}_{0}|$. This frequency entrainment is yet another signature of the dressed microwave resonator behaving as a self-sustained oscillator [2, 3].

### v2

@@ II.2 Effective photon action to fourth order in $g$
Figure 2: Real and imaginary parts of the retarded, and imaginary part of the Keldysh photon self energy plotted in panels (a, c, e) against DQD detuning for $V=10\omega_{0}$, and in panels (b, d, f) against bias voltage for $\varepsilon=(\sqrt{3}/2)\omega_{0}$. In both cases our $L/R$-symmetric parameter set implies that $\Pi^{\text{R}}_{LL}=\Pi^{\text{R}}_{RR}$ and $\Pi^{\text{R}}_{LR}=\Pi^{\text{R}}_{RL}.$ The dashed gridline in panel (e) shows the resonance condition, $\varepsilon=(\sqrt{3}/2)\omega_{0}$, and in (f) the threshold condition for the bias voltage, $V=\omega_{0}.$ The width of the transition region around the dashed gridline is set by $\Gamma$.
@@ III.2 Fluctuations: Fokker-Planck and beyond
Figure 4: Limit cycle radius, $r_{\ast}$, determined respectively by the saddle-point equations, $r_{\text{sp}}$, or the value at which the Wigner function attains its maximum, $r_{\text{W}}$, and plotted together with the root mean square of $r$, calculated from the Wigner function. The inset shows a zoom in on the bifurcation region, $g\sim g_{c}$, revealing a slight difference in the predicted values for the critical coupling, $g_{c}$ (gridline marks $\tilde{\kappa}=0$).
@@
for a distribution function, $P(\bar{\phi},\phi,t)$ and with diffusion constant $D_{0}$. Transforming to polar representation, $\phi=re^{i\theta}$, and employing Wirtinger derivatives, $\partial_{\phi}=e^{-i\theta}(\partial_{r}-ir^{-1}\partial_{\theta})/2$ and $\partial_{\bar{\phi}}=e^{i\theta}(\partial_{r}+ir^{-1}\partial_{\theta})/2$, this transforms to
@@
where $c$ ensures normalization, $\int_{0}^{\infty}\!dr\,2\pi rP_{0}(r)=1$, and the subscript 0 indicates that $\Lambda_{2,3,4,5}$ are all omitted. The same solution was found in Ref. [12], for a LME with jump operators representing 1-photon drive and 2-photon losses. Without the nonlinearity, one can relate the distribution function to the Wigner function as $P(\bar{\phi},\phi,t)=W(\bar{\phi}/\sqrt{2},\phi/\sqrt{2},t)/2$. In the general nonlinear case, however, this identification is not strictly valid. In particular, we note that the solution (27) is not a true Wigner function in the sense that the corresponding density matrix is not strictly positive semi-definite.
The solution (27) is valid only if $\Lambda_{1}^{\prime\prime}/D_{0}>0$, which is the case for $V>\omega_{0}$ (cf. Fig. 3), in which case the Hopf bifurcation takes place as $\tilde{\kappa}$ passes through zero. Below and above the bifurcation, the maximum of the probability distribution is attained, respectively, at $r=0$, or $r=r_{\ast}=\sqrt{-\tilde{\kappa}/\Lambda_{1}^{\prime\prime}}$, consistent with the saddle point solution found above. This bifurcation is shown in Fig. 4 as a sharp transition from the trivial fixed point, $r=0$, to the limit cycle at $r=r_{\ast}$ at a critical coupling, $g_{c}$, close to and above which $r_{\ast}\sim|g-g_{c}|^{1/2}$. The orange curve is determined as the fixed point of the classical saddle point equations, corresponding to the $r$ at which $P_{0}(r)$ reaches its maximum. The dark blue curve shows the exact radius of the limit cycle obtained as the radius at which the Wigner function attains its maximum. The Wigner functions is calculated using the LME, which treats the interaction between photons and electrons exactly. From the inset, the true bifurcation is seen to take place for a slightly smaller coupling than the value $g_{c}=0.043\omega_{0}$, determined from the saddle-point equations by the vanishing of $\tilde{\kappa}=0$. Fig. 4 also shows that the saddle-point value for the limit-cycle radius, $r_{\text{sp}}$, quickly becomes inaccurate for $g/\omega_{0}\gtrsim 0.06.$
@@
We solve this eigenvalue problem numerically  [^1] The calculations were performed using the Mathematica function NDeigensystem, where the area of the cells in the discrete spatial grid can be specified., and find that except for the zero eigenvalue corresponding to the steady state, all eigenvalues of the resulting differential operator $\mathcal{L}_{1}+\mathcal{L}_{2}$ are real and negative.
Fig. 5 shows a comparison of the steady state solution, $P$, to the previous solution, $P_{0}$, obtained from Eq. (27), and to the exact Wigner function, $W$, obtained from the LME. While all solutions preserve the bifurcation, Fig. 5(d) shows that the inclusion of $\Lambda_{2}$ and $\Lambda_{5}$ lowers $g_{c}$ towards the correct value obtained from the Wigner function. Calculating the photon numbers as
[…]
the three different distributions, $P_{i}=\{P_{0},P,P_{W}\}$, with $P_{W}=W(\bar{\phi}/\sqrt{2},\phi/\sqrt{2},t)/2$, shown in Figs. 5(d) and (f), correspond to photon numbers $N_{\rm ph}=\{2.99,2.85,3.07\}$ and $N_{\rm ph}=\{4.54,4.48,6.34\}$, respectively. Despite their different shapes, both $P_{0}$ and $P$ agree very well on this lowest moment, which however disagrees markedly with that of the Wigner function already for $g=0.05$. Note that for the parameters considered here, the linear shift performed earlier, $\delta X^{\rm cl}=2gn_{R}/\omega_{0}$, is of the order of $10^{-2}$ and therefore is insignificant in these plots.
Figure 5: Wigner function (blue) overlaid by the limit-cycle radius, $r_{\ast}$, determined from the saddle-point equations (orange) [(a), (c), (e)] and radial probability distribution [(b), (d), (f)] obtained by solving the FP equation without/with ($P_{0}/P$) the nonlinear terms ($\Lambda_{2/5}$), plotted alongside $P_{W}=W(\bar{\phi}/\sqrt{2},\phi/\sqrt{2},t)/2$ with $\phi=re^{i\theta}$. From top to bottom the coupling strength increases as $g/\omega_{0}=[0.01,0.042,0.05]$. The dashed gridline in panel (f) marks the maximum of $P_{0}$, i.e. the saddle-point value of the limit-cycle radius, $r_{\ast}$.
@@ III.3 Entrainment
we solve for $(X(t),P(t))=({\rm{Re}}[\phi],{\rm{Im}}[\phi])\sqrt{2}$. The corresponding phase-space trajectories are shown in Fig.6 together with the Lissajous curves traced out by $(X(t),F(t))$, revealing frequency entrainment when the amplitude of the force exceeds a critical strength, found here to be $A_{c}=0.044$. The rightmost column shows the power spectral density (PSD), calculated from the auto-correlation function of $X(t)$ (cf. Appendix B), which displays a clear transfer of spectral weight from $\tilde{\omega}_{0}$ to $\Omega$ as the amplitude of the force is increased. Note that the entrainment can also be found below the bifurcation point, $g<g_{c}.$ In this case any finite driving amplitude makes it possible to sustain a closed phase-space trajectory with finite radius, despite the fixed point of the undriven system being $r=0$ (see Fig. 6(d)). The radius of the trajectory increases with increasing $A,$ and decreasing detuning, $|\Omega-\tilde{\omega}_{0}|$. This frequency entrainment is yet another signature of the dressed microwave resonator behaving as a self-sustained oscillator [2, 3].